i) give a recursive definition of the set P of all positive integers greater than 0
II) formulate tha appropriate induction principle
III) use mathematical induction to prove that
11+15+19 +... + (4n+7) = 2n^2 + 9n for all positive n>0

Answer by jibirajeev(1)

My definition p(n)=p(n-1)+1
11+15+19..+(4n+7)+(4(n+1)+7)=2(n+1)^2+9(n+1)
=2(n^2+2n+1)+9n+1
= 2n^2+9n + 4n+2+9
=2n^2+9n + 4n+11
Therefore works for the step up case
= 4(n+1)+7=4n+11